# A Taxonomy of Multi-Layered Runtime Guardrails for Designing Foundation Model-Based Agents: Swiss Cheese Model for AI Safety by Design

### v2

@@ V-C Threats to Validity
We prepared a guardrails taxonomy and conducted a comparative analysis of its components to help the reader better understand their design and evaluation. We critically examined the strength and consistency of relationships in the selected studies for a reliable taxonomy structure and drew conclusions. Nonetheless, the generalizability of guardrails to different contexts and types in FM-based agents remains a potential limitation, and specific adaptations might be necessary for certain systems, such as those used in healthcare or financial organizations. Additionally, our findings face a challenge as some of them may become outdated due to the rapid evolution of FM-based agents and associated guardrails.
VI Conclusion and Future Work
To better understand runtime guardrails in FM-based agents, this paper developed and presented a comprehensive guardrail taxonomy and provided a comparative study of its components based on a systematic literature review guided by the Swiss Cheese Model for AI safety. Our proposed taxonomy categorizes guardrails based on their fundamental objectives, essential quality attributes, performed actions, targeted aspects, covered scopes, employed rules, autonomy, modalities, and underlying techniques.
Our key findings emphasize that the effective integration of guardrails mitigates risk such as biases, harmful content, and unintended behaviors in FM-based agents. Practitioners can use the proposed taxonomy to improve the design of guardrails to ensure the behavior and decisions of FM-based agents are responsible and safe. Moreover, the taxonomy provides a theoretical framework that researchers and policymakers can use to develop guidelines and standards that promote AI-safety-by-design in FM-based agents.
In the future, we plan to develop a multilayer architecture for guardrails in FM-based agents, integrating various design options outlined in this taxonomy. Additionally, we aim to evaluate the effectiveness of multi-layered guardrails in real-world agent deployments, further refining the taxonomy and contributing to best practices in AI safety.

### v3

@@ VI Threats to Validity
We prepared a guardrails taxonomy and conducted a comparative analysis of its components to help the reader better understand their design and evaluation. We critically examined the strength and consistency of relationships in the selected studies to develop a reliable taxonomy and reference architecture for designing built-in multi-layered runtime guardrails. Finally, we draw conclusions. Nonetheless, the generalizability of guardrails to different contexts and types in FM-based agents remains a potential limitation. Specific adaptations might be necessary for certain systems, such as those used in healthcare or financial organizations.
VII Conclusion and Future Work
To advance the understanding of runtime guardrail design in FM-based agents, this paper presents a comprehensive taxonomy of guardrail design based on the results of an SLR. Our taxonomy categorizes guardrails based on their essential quality attributes and key design dimensions, including guardrail actions and targets, employed rules, guardrail sources, modality, and underlying models. Building on this taxonomy, we propose a novel Swiss Cheese Model for AI safety - a reference architecture for designing built-in, multi-layered guardrails in FM-based agents, which includes three dimensions: quality attributes, pipelines, and artifacts. In the future, we plan to develop guardrail services for a scientific agent platform, implementing the proposed reference architecture and integrating various design options outlined in the taxonomy.
Appendix A: List of Selected Studies
  • SS [1] M. Pawagi, V. Kumar, Guardrails: Automated suggestions for clarifying ambiguous purpose statements, in: Proceedings of the 16th Annual ACM India Compute Conference, COMPUTE ’23, Association for Computing Machinery, New York, NY, USA, 2023, p. 55–60. doi:10.1145/3627217.3627234.
  • SS [2] A. Khorramrouz, S. Dutta, A. Dutta, A. R. KhudaBukhsh, Down the toxicity rabbit hole: Investigating palm 2 guardrails, arXiv preprint arXiv:2309.06415 (2023). doi:https://doi.org/10.48550/arXiv.2309.06415.
  • SS [3] Y. Dong, R. Mu, G. Jin, Y. Qi, J. Hu, X. Zhao, J. Meng, W. Ruan, X. Huang, Building guardrails for large language models, arXiv preprint arXiv:2402.01822 (2024). doi:https://doi.org/10.48550/arXiv.2402.01822.
  • SS [4] N. Mangaokar, A. Hooda, J. Choi, S. Chandrashekaran, K. Fawaz, S. Jha, A. Prakash, Prp: Propagating universal perturbations to attack large language model guard-rails, arXiv preprint arXiv:2402.15911 (2024). doi:https://doi.org/10.48550/arXiv.2402.15911.
  • SS [5] M. Anderljung, J. Barnhart, J. Leung, A. Korinek, C. O’Keefe, J. Whittlestone, S. Avin, M. Brundage, J. Bullock, D. Cass-Beggs, et al., Frontier ai regulation: Managing emerging risks to public safety, arXiv preprint arXiv:2307.03718 (2023). doi:https://doi.org/10.48550/arXiv.2307.03718.
  • SS [6] M. Liffiton, B. E. Sheese, J. Savelka, P. Denny, Codehelp: Using large language models with guardrails for scalable support in programming classes, in: Proceedings of the 23rd Koli Calling International Conference on Computing Education Research, Koli Calling ’23, Association for Computing Machinery, New York, NY, USA, 2024, pp. 1–11. doi:10.1145/3631802.3631830.
  • SS [7] Z. Zhang, Y. Lu, J. Ma, D. Zhang, R. Li, P. Ke, H. Sun, L. Sha, Z. Sui, H. Wang, et al., Shieldlm: Empowering llms as aligned, customizable and explainable safety detectors, arXiv preprint arXiv:2402.16444 (2024). doi:https://doi.org/10.48550/arXiv.2402.16444.
  • SS [8] T. Rebedea, R. Dinu, M. N. Sreedhar, C. Parisien, J. Cohen, NeMo guardrails: A toolkit for controllable and safe LLM applications with programmable rails, in: Y. Feng, E. Lefever (Eds.), Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing: System Demonstrations, Association for Computational Linguistics, Singapore, 2023, pp. 431–445. doi:10.18653/v1/2023.emnlp-demo.40.
  • SS [9] Y. Wang, L. Singh, Adding guardrails to advanced chatbots, arXiv preprint arXiv:2306.07500 (2023). doi:https://doi.org/10.48550/arXiv.2306.07500.
  • SS [10] M. Shanahan, Talking about large language models, Commun. ACM 67 (2) (2024) 68–79. doi:10.1145/3624724.
  • SS [11] W. Du, Q. Li, J. Zhou, X. Ding, X. Wang, Z. Zhou, J. Liu, Finguard: A multimodal aigc guardrail in financial scenarios, in: Proceedings of the 5th ACM International Conference on Multimedia in Asia, MMAsia ’23, Association for Computing Machinery, New York, NY, USA, 2024, pp. 1–3. doi:10.1145/3595916.3626351.
  • SS [12] A. Wei, N. Haghtalab, J. Steinhardt, Jailbroken: How does llm safety training fail?, Advances in Neural Information Processing Systems 36 (2024). doi:10.48550/arXiv.2307.02483.
  • SS [13] J. Zhao, K. Chen, X. Yuan, Y. Qi, W. Zhang, N. Yu, Silent guardian: Protecting text from malicious exploitation by large language models, arXiv preprint arXiv:2312.09669 (2023). doi:https://doi.org/10.48550/arXiv.2312.09669.
  • SS [14] X. Qi, Y. Zeng, T. Xie, P.-Y. Chen, R. Jia, P. Mittal, P. Henderson, Fine-tuning aligned language models compromises safety, even when users do not intend to!, arXiv preprint arXiv:2310.03693 (2023). doi:https://doi.org/10.48550/arXiv.2310.03693.
  • SS [15] J. Mökander, J. Schuett, H. R. Kirk, L. Floridi, Auditing large language models: a three-layered approach, AI and Ethics (2023) 1–31doi:https://doi.org/10.1007/s43681-023-00289-2.
  • SS [16] S. Banerjee, S. Layek, R. Hazra, A. Mukherjee, How (un) ethical are instruction-centric responses of llms? unveiling the vulnerabilities of safety guardrails to harmful queries, arXiv preprint arXiv:2402.15302 (2024). doi:https://doi.org/10.48550/arXiv.2402.15302.
  • SS [17] S. Ee—Researcher, Z. W. Director, Adapting cybersecurity frameworks to manage frontier AI risks, Institute for AI Policy and Strategy (IAPS) (2023).
    URL https://static1.squarespace.com/static/64edf8e7f2b10d716b5ba0e1/t/6528c5c7f912f74fbd03fc34/1697170896984/Adapting+cybersecurity+frameworks+to+manage+frontier+AI+risks.pdf
  • SS [18] P. Rai, S. Sood, V. K. Madisetti, A. Bahga, Guardian: A multi-tiered defense architecture for thwarting prompt injection attacks on llms, Journal of Software Engineering and Applications 17 (1) (2024) 43–68. doi:https://doi.org/10.4236/jsea.2024.171003.
  • SS [19] A. Kumar, S. Singh, S. V. Murty, S. Ragupathy, The ethics of interaction: Mitigating security threats in llms, arXiv preprint arXiv:2401.12273 (2024). doi:https://doi.org/10.48550/arXiv.2401.12273.
  • SS [20] X. Shen, Z. Chen, M. Backes, Y. Shen, Y. Zhang, ”do anything now”: Characterizing and evaluating in-the-wild jailbreak prompts on large language models, arXiv preprint arXiv:2308.03825 (2023). doi:https://doi.org/10.48550/arXiv.2308.03825.
  • SS [21] A. Kumar, C. Agarwal, S. Srinivas, S. Feizi, H. Lakkaraju, Certifying llm safety against adversarial prompting, arXiv preprint arXiv:2309.02705 (2023). doi:https://doi.org/10.48550/arXiv.2309.02705.
  • SS [22] Y. Zeng, H. Lin, J. Zhang, D. Yang, R. Jia, W. Shi, How johnny can persuade llms to jailbreak them: Rethinking persuasion to challenge ai safety by humanizing llms, arXiv preprint arXiv:2401.06373 (2024). doi:https://doi.org/10.48550/arXiv.2401.06373.
  • SS [23] B. Wei, K. Huang, Y. Huang, T. Xie, X. Qi, M. Xia, P. Mittal, M. Wang, P. Henderson, Assessing the brittleness of safety alignment via pruning and low-rank modifications, arXiv preprint arXiv:2402.05162 (2024). doi:https://doi.org/10.48550/arXiv.2402.05162.
  • SS [24] S. Goyal, M. Hira, S. Mishra, S. Goyal, A. Goel, N. Dadu, D. Kirushikesh, S. Mehta, N. Madaan, Llmguard: Guarding against unsafe llm behavior, in: Proceedings of the AAAI Conference on Artificial Intelligence, Vol. 38(21), 2024, pp. 23790–23792. doi:https://doi.org/10.1609/aaai.v38i21.30566.
  • SS [25] R. R. Llaca, V. Leskoschek, V. C. Paiva, C. Lupău, P. Lippmann, J. Yang, Student-teacher prompting for red teaming to improve guardrails, in: Proceedings of the ART of Safety: Workshop on Adversarial testing and Red-Teaming for generative AI, 2023, pp. 11–23. doi:http://dx.doi.org/10.18653/v1/2023.artofsafety-1.2.
  • SS [26] Z. Wang, F. Yang, L. Wang, P. Zhao, H. Wang, L. Chen, Q. Lin, K.-F. Wong, SELF-GUARD: Empower the LLM to safeguard itself, in: K. Duh, H. Gomez, S. Bethard (Eds.), Proceedings of the 2024 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies (Volume 1: Long Papers), Association for Computational Linguistics, Mexico City, Mexico, 2024, pp. 1648–1668. URL https://aclanthology.org/2024.naacl-long.92
  • SS [27] B. Wang, W. Chen, H. Pei, C. Xie, M. Kang, C. Zhang, C. Xu, Z. Xiong, R. Dutta, R. Schaeffer, et al., Decodingtrust: A comprehensive assessment of trustworthiness in gpt models, in: Advances in Neural Information Processing Systems, 2023, pp. 1–110. doi:https://doi.org/10.48550/arXiv.2306.11698.
  • SS [28] R. Bommasani, D. A. Hudson, E. Adeli, R. B. Altman, S. Arora, S. von Arx, M. S. Bernstein, J. Bohg, A. Bosselut, E. Brunskill, E. Brynjolfsson, S. Buch, D. Card, R. Castellon, N. S. Chatterji, A. S. Chen, K. Creel, J. Q. Davis, D. Demszky, C. Donahue, M. Doumbouya, E. Durmus, S. Ermon, J. Etchemendy, K. Ethayarajh, L. Fei-Fei, C. Finn, T. Gale, L. Gillespie, K. Goel, N. D. Goodman, S. Grossman, N. Guha, T. Hashimoto, P. Henderson, J. Hewitt, D. E. Ho, J. Hong, K. Hsu, J. Huang, T. Icard, S. Jain, D. Jurafsky, P. Kalluri, S. Karamcheti, G. Keeling, F. Khani, O. Khattab, P. W. Koh, M. S. Krass, R. Krishna, R. Kuditipudi, et al., On the opportunities and risks of foundation models, CoRR abs/2108.07258 (2021). arXiv:2108.07258, doi:https://doi.org/10.48550/arXiv.2108.07258.
  • SS [29] L. Weidinger, M. Rauh, N. Marchal, A. Manzini, L. A. Hendricks, J. Mateos-Garcia, S. Bergman, J. Kay, C. Griffin, B. Bariach, et al., Sociotechnical safety evaluation of generative ai systems, arXiv preprint arXiv:2310.11986 (2023). doi:https://doi.org/10.48550/arXiv.2310.11986.
  • SS [30] D. Kang, D. Raghavan, P. Bailis, M. Zaharia, Model assertions for monitoring and improving ml models, Proceedings of Machine Learning and Systems 2 (2020) 481–496. doi:https://doi.org/10.48550/arXiv.2003.01668.
  • SS [31] Z. Yuan, Z. Xiong, Y. Zeng, N. Yu, R. Jia, D. Song, B. Li, Rigorllm: Resilient guardrails for large language models against undesired content, arXiv preprint arXiv:2403.13031 (2024). doi:https://doi.org/10.48550/arXiv.2403.13031.
  • SS [32] Z. Chu, Y. Wang, L. Li, Z. Wang, Z. Qin, K. Ren, A causal explainable guardrails for large language models, arXiv preprint arXiv:2405.04160 (2024). doi:https://doi.org/10.48550/arXiv.2405.04160
Appendix B: QAC Scores for the Selected Studies
Selected Study (SS) No. QAC1 QAC2 QAC3 QAC4 QAC5
SS-1 2 2 2 2 2
SS-2 4 5 5 5 4
SS-3 4 3 2 3 2
SS-4 3 3 3 3 4
SS-5 2 4 3 3 5
SS-6 1 5 4 3 5
SS-7 5 4 4 4 3
SS-8 5 4 4 5 3
SS-9 0 1 2 4 3
SS-10 0 2 2 3 4
SS-11 3 3 1 3 2
SS-12 3 2 3 3 2
SS-13 4 3 3 4 3
SS-14 5 4 4 5 4
SS-15 5 4 3 4 5
SS-16 5 4 4 5 4
SS-17 5 4 3 5 4
SS-18 5 4 4 5 5
SS-19 4 3 2 4 3
SS-20 5 4 4 5 4
SS-21 5 4 4 4 5
SS-22 5 5 4 5 4
SS-23 5 4 4 5 4
SS-24 5 4 2 2 3
SS-25 5 4 4 4 5
SS-26 5 4 4 5 4
SS-27 5 4 5 5 4
SS-28 5 4 4 3 4
SS-29 4 4 4 3 4
SS-30 5 4 4 4 5
SS-31 5 5 4 4 5
SS-32 5 4 4 4 5